# Split Two-Periodic Aztec Diamond

### v3

@@ 4.3.2. Asymptotics on the $\alpha$-side when $\beta>\alpha$
(84) $\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{w,\uparrow}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]\\ =-C_{1}\exp\big{[}\varphi_{\alpha,2}(z_{2}^{*};x_{2}-x_{1},y_{2}-y_{1})\big{]}N^{-1}\int_{\mathbb{R}}\text{d}r\int_{\textrm{i}\mathbb{R}}\text{d}s\exp\Big{[}\psi_{2}^{\prime\prime}(z_{2}^{*})(r^{2}-s^{2})\Big{]}+o(N^{-1})$
By Lemma 4.2, we know that $\operatorname{Re}\psi^{\prime\prime}_{2}(z_{2}^{*})<0$, since $z_{2}^{*}$ is a maximum of $\operatorname{Re}\psi^{\prime\prime}_{2}(z)$, and so the double integral is convergent. We repeat the procedure with the contour $\gamma_{w,\downarrow}$, noting that
(85) $\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)F_{\alpha,2}(z)}{z(1+2g_{\alpha,\beta}(w))}=\overline{C}_{1}N^{-1/2}(s-r)+O(N^{-1})$
since the contour has been deformed from below the branch cut. Thus we get,
(86) $\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{w,\downarrow}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]\\ =\overline{C}_{1}\exp\big{[}\varphi_{\alpha,2}(z_{2}^{*};x_{2}-x_{1},y_{2}-y_{1})\big{]}N^{-1}\int_{\mathbb{R}}\text{d}r\int_{\textrm{i}\mathbb{R}}\text{d}s\exp\Big{[}\psi_{2}^{\prime\prime}(z_{2}^{*})(r^{2}-s^{2})\Big{]}+o(N^{-1})$
Combining equations (84) and (86), we conclude that for $(x,y)\in\mathfrak{C}$,
@@
where $i,j\in\{0,1\}$. At the beginning of the proof, we assumed $z_{2}^{*}\in[-\beta^{2},-\alpha^{2})$. If instead $z_{2}^{*}\in(-\alpha^{-2},-\beta^{-2}]$, the procedure would be largely the same as above. The only difference being that we would split up the contour $\gamma_{w,2}$ as it would be the one that intersects $\gamma_{s,2}$. ∎
4.4. Asymptotics of $I^{\alpha}_{2,1}$
The asymptotics of this double contour integral does depend on the region of the model, so we will consider the asymptotics in the rough and smooth regions separately. Before we do so, we will repeat a procedure seen in the proof of Corollary 2.1 in Section 3.4 on the double contour integral $I^{\alpha}_{2,1}$. Namely, we will deform the $w$-contours through the branch cuts and onto the other sheet of the Riemann surface. Doing so we obtain,
(87) $\Big{[}I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}\\ \times r_{\alpha,1}(z)^{m-\frac{N}{2}}\frac{2F_{\alpha,1}(w)F_{\beta,2}(w)F_{\alpha,2}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,1}(z)$
4.4.1. Asymptotics of $I^{\alpha}_{2,1}$ in the Rough Region
Here we will assume $(x,y)\in\mathfrak{R}$ and $x<\nicefrac{{1}}{{2}}$ and prove part (1) of Proposition 2.3.
Proof.
We start by deforming the contours to the contours of steepest descent and ascent detailed in Section 4.1 and Lemma 4.4 and depicted in Figure 11. This deformation produces a single contour integral due to the residue at $z=w$, but the contour integral amounts to zero since it contains the matrix product $F_{\alpha,2}(z)F_{\alpha,1}(z)=0$. We are left with the following double contour integral, which we write in terms of the saddle functions.
@@ 4.4.1. Asymptotics of $I^{\alpha}_{2,1}$ in the Rough Region
(88) $\oint_{\gamma_{R,D}}\oint_{\gamma_{R,A}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,1}(z)^{m-\frac{N}{2}}\frac{2F_{\alpha,1}(w)F_{\beta,2}(w)F_{\alpha,2}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,1}(z)\\ =\oint_{\gamma_{R,D}}\oint_{\gamma_{R,A}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,1}(w)F_{\beta,2}(w)F_{\alpha,2}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,1}(z)\\ \times\exp\left[N\left(\psi_{1}(w;x,y)-\psi_{1}(z;x,y)\right)+\varphi_{\alpha,1}(w;x_{2},y_{2})-\varphi_{\alpha,1}(z;x_{1},y_{1})\right]\$
Now we can use saddle point methods to compute the leading order decay of this integral. Since there are two different saddle points, we must consider the contributions from the four combinations. [^11] We can let $w=z_{3}^{*}$ or $w=z_{4}^{*}$. The same applies for $z$, thus resulting in four combinations. The contributions where different saddle points are used for $w$ and $z$ are immediate. We get the leading order term
(89) $\frac{F_{\alpha,1}(z_{l}^{*})F_{\beta,2}(z_{l}^{*})F_{\alpha,2}(z_{l}^{*})F_{\alpha,1}(z_{k}^{*})\exp\Big{[}N(\psi_{1}(z_{l}^{*})-\psi_{1}(z_{k}^{*}))+\varphi_{1}(z_{l}^{*})-\varphi_{1}(z_{k}^{*})+2\textrm{i}\,\theta_{lk}(x,y)\Big{]}}{\pi z_{k}^{*}(z_{k}^{*}-z_{l}^{*})(1+2g_{\alpha,\beta}(z_{l}^{*}))N\sqrt{|\psi^{\prime\prime}(z_{k}^{*})\psi^{\prime\prime}(z_{l}^{*})|}}$
where either $l=3$ and $k=4$ or $l=4$ and $k=3$. The function $\theta_{lk}(x,y)$ depends on the angle of the steepest ascent and descent contours around the saddle points. The term $\psi_{1}(z_{l}^{*})-\psi_{1}(z_{k}^{*})$ is purely imaginary, and thus does not contribute to the decay of the term. The absolute value of the above can be bounded by $CN^{-1}$, where $C$ is some constant matrix.
Now we want to consider the leading order contributions from the terms where we consider $z$ and $w$ at the same saddle point. We start by expanding the function $\psi_{1}(z)$ around $z=z_{k}^{*}$,
$\psi_{1}(z)=\psi_{1}(z_{k}^{*})+\psi_{1}^{\prime\prime}(z_{k}^{*})(z-z_{k}^{*})^{2}+O((z-z_{k}^{*})^{3})$
In addition, we make the following change of variables,
(90) $z-z_{k}^{*}=sN^{-1/2}\qquad\qquad w-z_{k}^{*}=rN^{-1/2}$
Combining this gives,
(91) $N\Big{[}\psi_{1}(w)-\psi_{1}(z)\Big{]}=\psi_{1}^{\prime\prime}(z_{k}^{*})(s^{2}-r^{2})+O(N^{-1/2})$
We also apply Lemma 4.5 to get,
(92) $\frac{2F_{\alpha,1}(z)F_{\beta,2}(z)F_{\alpha,2}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}=\frac{2F_{\alpha,1}(z_{k}^{*})F_{\beta,2}(z_{k}^{*})F_{\alpha,2}(z_{k}^{*})F^{\prime}_{\alpha,1}(z_{k}^{*})}{1+2g_{\alpha,\beta}(z_{k}^{*})}N^{-1/2}(s-r)+O(N^{-1})$
Combining all this we get the leading order asymptotics,
(93) $\frac{1}{(2\pi)^{2}}\frac{2F_{\alpha,1}(z_{k}^{*})F_{\beta,2}(z_{k}^{*})F_{\alpha,2}(z_{k}^{*})F^{\prime}_{\alpha,1}(z_{k}^{*})}{z_{k}^{*}(1+2g_{\alpha,\beta}(z_{k}^{*}))}N^{-1}\int_{\Gamma_{r}}\int_{\Gamma_{s}}\exp\big{[}\psi_{1}^{\prime\prime}(z_{k}^{*})(s^{2}-r^{2})\big{]}\text{d}s\text{d}r$
Where $\Gamma_{r}$ and $\Gamma_{s}$ are lines through the origin whose angles depend on the angle of the contours of steepest descent and ascent, respectively. Combining the results from equations 89, 93 and Lemma LABEL:lem:decaypart we obtain the results presented in part (1) of Proposition 2.3. ∎
4.4.2. Asymptotics of $I^{\alpha}_{2,1}$ in the Smooth Region
The results in parts (2) and (3) of Proposition 2.3 are proven using the same procedure as the proof of Proposition 2.2, located in Section 4.3. The main difference we need to address is that $\psi_{1}(z)$ has three saddle points on the interval $[-\alpha^{-2},-\alpha^{2}]$, where $\psi_{2}(z)$ only had one. However, Lemma 4.3 gives all the necessary facts needed to repeat the procedure given in Section 4.3. In particular, we are able to deform the contours in the same way described in Section 4.3.
Proof.
We deform the $z$-contour through the central saddle point, $z_{1}^{*}$, which is a local maximum of $\operatorname{Re}\psi_{1}(z)$ along the interval (see Lemma 4.2). We see from Lemma 4.3, that we can utilize the same placement of contours as we did in the proof of Proposition 2.2. Thus, if $(x,y)\not\in\mathfrak{C}$ we can conclude that
$|I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq c_{2}e^{-c_{1}N}$
for $i,j\in\{0,1\}$ and some positive constants $c_{1}$ and $c_{2}$. If $(x,y)\in\mathfrak{C}$ we have,
(94) $\Big{[}I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=(\pm C_{2}\mp\overline{C}_{2})\exp\big{[}\varphi_{\alpha,2}(z_{1}^{*};x_{2}-x_{1},y_{2}-y_{1})\big{]}N^{-1}\\ \times\int_{\mathbb{R}}\text{d}r\int_{\textrm{i}\mathbb{R}}\text{d}s\exp\Big{[}\psi_{2}^{\prime\prime}(z_{1}^{*})(r^{2}-s^{2})\Big{]}+o(N^{-1})$
where
(95) $C_{2}=\frac{2F_{\alpha,2}(z_{1}^{*})F_{\beta,1}(z_{1}^{*})F_{\alpha,1}(z_{1}^{*})F^{\prime}_{\alpha,2}(z_{1}^{*})}{1+2g_{\alpha,\beta}(z_{1}^{*})}$
The plus or minus signs in equation (94) depend on whether $z_{1}^{*}>-\alpha^{2}$ or $z_{1}^{*}<-\alpha^{-2}$. ∎

### v4

@@ 4.3.1. Case 1: The asymptotic coordinate is in the rough region
(86) $\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br},\uparrow}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]\\ =-C_{1}\exp\big{[}\varphi_{\alpha,2}(z_{2}^{*};x_{2}-x_{1},y_{2}-y_{1})\big{]}N^{-1}\int_{\mathbb{R}}\text{d}r\int_{\textrm{i}\mathbb{R}}\text{d}s\exp\Big{[}\psi_{2}^{\prime\prime}(z_{2}^{*})(r^{2}-s^{2})\Big{]}+o(N^{-1})$
By Lemma 4.2, we know that $\operatorname{Re}\psi^{\prime\prime}_{2}(z_{2}^{*})<0$, since $z_{2}^{*}$ is a maximum of $\operatorname{Re}\psi^{\prime\prime}_{2}(z)$, and so the integral above is convergent. We repeat the procedure with the contour $\gamma_{\text{Br},\downarrow}$, noting that
(87) $\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)F_{\alpha,2}(z)}{z(1+2g_{\alpha,\beta}(w))}=C_{2}N^{-1/2}(s-r)+O(N^{-1})$
where,
(88) $C_{2}=\frac{2F_{\alpha,2}(z_{2}^{*})\overline{F_{\beta,1}}(z_{2}^{*})F_{\alpha,1}(z_{2}^{*})F^{\prime}_{\alpha,2}(z_{2}^{*})}{z_{2}^{*}(1+2\overline{g_{\alpha,\beta}}(z_{2}^{*}))}$
since we are approaching the branch cut from below. Thus we get,
(89) $\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br},\downarrow}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]\\ =C_{2}\exp\big{[}\varphi_{\alpha,2}(z_{2}^{*};x_{2}-x_{1},y_{2}-y_{1})\big{]}N^{-1}\int_{\mathbb{R}}\text{d}r\int_{\textrm{i}\mathbb{R}}\text{d}s\exp\Big{[}\psi_{2}^{\prime\prime}(z_{2}^{*})(r^{2}-s^{2})\Big{]}+o(N^{-1})$
Combining equations (86) and (89), we conclude that for $(x,y)\in\mathfrak{S}\cap\mathfrak{C}$,
@@
where $i,j\in\{0,1\}$. ∎
4.4. Local Asymptotics of $I^{\alpha}_{2,1}$
The asymptotics of the contour integral $I^{\alpha}_{2,1}$ are much simpler, in the sense that they only depend on $(x,y)$ being away from any macroscopic boundary and $x<\nicefrac{{1}}{{2}}$. We use the work done in Section 4.2 as a launching off point. In Lemma 4.5 we showed that the single integral that arises from manipulating the contours decays exponentially in $N$, so to prove Proposition 2.2 we only need to show that the following integral decays exponentially as well,
@@ 4.4. Local Asymptotics of $I^{\alpha}_{2,1}$
(90) $\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)F_{\varepsilon,1}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{1}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,1}(z;x_{1},y_{1})\right]$
From Lemma 4.2, we know that $\psi_{1}(z)$ has a maximum at $z_{1}^{*}$, so $\gamma_{0,1}$ should be deformed to the steepest ascent contour through this point $\gamma_{s,1}$. After this deformation, we only need to show that
(91) $\operatorname{Re}\psi_{2}(w)<\operatorname{Re}\psi_{1}(z_{1}^{*})$
for all $w\in\gamma_{\text{Br}}$. Once this point is justified, the proof will follow the proof of case (2) in Proposition 2.1. In particular, adding the statement in Lemma 4.1 to the statement above is enough to conclude
$\exp\Big{[}N\big{(}\psi_{2}(w;x,y)-\psi_{1}(z;x,y)\big{)}\Big{]}$
is $O(e^{-cN})$ for some $c>0$, uniformly for $(z,w)\in\gamma_{z}\times\gamma_{\text{Br},1}\cup\gamma_{\text{Br},2}$.
We will now prove the claim in equation (91). We cannot immediately apply Corollary 4.1 as we did in Section 4.3.2, since now we are trying to compare $\psi_{1}(z_{1}^{*})$ to $\psi_{2}(w)$. We use the fact that $r_{\alpha,2}(w)=r_{\alpha,1}(w)^{-1}$ and the fact that $x<\nicefrac{{1}}{{2}}$ to write,
$\operatorname{Re}\psi_{2}(w)$ $=(1+y)\log|w|-\log|w-1|+\left(\frac{1}{2}-x\right)\log|r_{\alpha,2}(w)|$
$=(1+y)\log|w|-\log|w-1|-\left(\frac{1}{2}-x\right)\log|r_{\alpha,1}(w)|$
(92) $\leq(1+y)\log|w|-\log|w-1|+\left(\frac{1}{2}-x\right)\log|r_{\alpha,1}(w)|=\operatorname{Re}\psi_{1}(w)$
where equality only occurs when $w\in(-\infty,-\alpha^{-2}]\cup[-\alpha^{2},0]$. We define the following sets,
$\Omega_{2,1}^{+}=\left\{w\in\mathbb{C}:\operatorname{Re}\psi_{2}(w)>\operatorname{Re}\psi_{1}(z_{1}^{*})\right\}$
$\Omega_{2,1}^{-}=\left\{w\in\mathbb{C}:\operatorname{Re}\psi_{2}(w)<\operatorname{Re}\psi_{1}(z_{1}^{*})\right\}$
We want to show that $\gamma_{\text{Br}}$ can be deformed such that $\gamma_{\text{Br}}\subset\Omega_{2,1}^{-}$. For $z_{1}^{*}\in(-\beta^{-2},-\beta^{2})$, this is immediate from combining Corollary 4.1 with the inequality in equation (92). Care should be taken in the case that $z_{1}^{*}\in(-\alpha^{-2},-\beta^{-2}]\cup[-\beta^{2},-\alpha^{2})$. In this case, we need the fact that the inequality in equation (92) is strict at $z_{1}^{*}$, so
$\operatorname{Re}\psi_{2}(z_{1}^{*})<\operatorname{Re}\psi_{1}(z_{1}^{*})$
When we combine this inequality with the continuity of $\operatorname{Re}\psi_{2}(w)$, we can conclude that $z_{1}^{*}$ is a positive distance away from $\Omega_{2,1}^{+}$, and so we can deform the contours in the desired manner.